OpenStudy (anonymous):

cos theta*cot(90-theta)=?

OpenStudy (campbell_st):

 \[\cot (90-\theta) = 1/\tan(90 - \theta)\] \[\tan (90 - \theta) = a/o\] then \[1/\tan(90 - \theta) = o/a\] so the problem now become \[\cos(\theta) \times \tan(\theta)\] and using the tan property \[\tan = \sin/\cos\] the problem is now \[\cos(\theta) \times \sin(\theta)/\cos(\theta) = \sin(\theta)\] Hope this helps